OpenStudy (anonymous):

The angle of minimum deviation for a prism of refractive index 3/2 is equal to angle of prism. What is the angle of the prism? ( cos 41 = 0.75) a) 31 degrees b) 41 degrees c) 62 degrees d) 82 degrees

OpenStudy (anonymous):

\[\mu=\sin ((A+\delta)/2)/\sin(A/2)\] here delta = A The RHS becomes sin(A)/sin(A/2)=2cos(A/2)=3/2 cos(A/2)=0.75 A/2=41 A=82.